# Position: A taxonomy for reporting and describing AI security incidents

### v1

@@ 3 Approach
To derive a taxonomy for reporting and describing AI security incidents, we followed a three step approach. Firstly, we reached out to twelve industry experts (working in AI red teaming, security industry, or AI governance), inquiring for their wish-list for a possible reporting scheme. We then compared the stated preferences with each other as well as existing incident collections or approaches similar to model cards. The derived, rudimentary scheme was then refined by iterative discussion rounds among four of the authors. When both draft and definitions had sufficiently converged, three of the authors used the tentative structure to tag incidents from existing databases. All appearing flaws were consecutively fixed, and the proposal was shared with the initial and additional experts for a broader, second round of feedback.
@@ 4 AI security incident taxonomy
Our taxonomy consists of five parts, where the last contains only an open text field for further comments and feedback, and is omitted from the following presentation. The parts are, (1) the incident’s meta data, (2) information about the targeted AI system, (3) specifics of the security incident, and lastly (4) implications of the incident. We now revisit the individual questions and references in each part. An example of the final questionnaire is depicted in Table 2-4 in the appendix.
4.1 Meta data (1.1-1.4)
The metadata is a short block containing the incident’s high-level information, and is, under variations, part of many databases (Huntr, ; McGregor, 2021; LVEs, ; ARVA, ; OECD, ). This information forms the first section with elements 1.1-1.4.
We opted to assign an incident ID, a name or ID of the incident reporter, their organization, and the alleged date (or interval, or unknown) of the incident. It is important to remark that an AI security incident may be reported anonymously. Although, as stated above, any question of our proposal can be dropped, we want to emphasize that the reporter and their organization may be pseudonymized or kept fully anonymous, depending on the specific usage of the taxonomy.
@@ 4.2 Targeted AI system
To understand an AI security incident, it is crucial to describe the AI system that formed part of the incident. Hence, in the second section of our taxonomy, we first collect information about the model, its life cycle stage and use, its origin, and application-specific information. Ideally, this information is directly imported from a model card or similar documentation to reduce required work.
@@
Model information (2.1-2.3). We first collect basic information about the model. This includes the model type, for example, whether a model is generative or predictive; a specific reply here could be a support vector machine. Such information was suggested by half of our experts and is used in several model card approaches (Mitchell et al., 2019; Arnold et al., 2019; Adkins et al., 2022). We then collect more specific information like the exact version of the model, information also encompassed in many model cards (Mitchell et al., 2019; Adkins et al., 2022) and requested by a fourth of our experts. Lastly, if possible, the model should be specified as detailed as possible (e.g., model weights), to enhance reproducibility, as suggested by another quarter of our experts.
@@
Life-cycle stage and intended use (2.4-2.7). In addition to the model within the AI system, understanding the system context is crucial. To this end, we also collect lifecycle information as specified by ISO [^9] ISO/IEC 22989 Information technology — Artificial intelligence — Artificial intelligence concepts and terminology, https://www.iso.org/standard/74296.html, as suggested by two of our experts. As suggested by the CSET AI Harm Framework (Hoffman & Frase, 2023) annotation guide, we also collect the domain where the AI system is deployed (Hoffmann et al., ). Intuitively, a deployment in physical domains has different implications than in the digital space. In addition, we keep track of the intended use of the AI system, loosely following the EU AI Act (Council of European Union, 2024), as done in many model card approaches (Mitchell et al., 2019; Hupont et al., 2024; Arnold et al., 2019; Adkins et al., 2022). To take into account specific exploits, we also advocate denoting a usage during the incident if different from the intended use.
@@
Model supply chain (2.8-2.10). Furthermore, it is crucial to understand possible (external) influences on the AI system stemming from the supply chain, as also suggested by two of our experts and in several AI model card approaches (Mitchell et al., 2019; Gebru et al., 2021; Adkins et al., 2022). Following scientific literature of AI security, we opted to divide the supply chain into three major areas: data (Mir, 2024), the model weights (Gu et al., 2019), and the code (Grosse et al., 2020; Shumailov et al., 2021). Any of these may be reused and thus allow the import of security vulnerabilities.
@@
Application (2.11). Instances of AI security incidents may need specific details to be reproducible or understandable. As an example, we put forward also collecting the system prompt in the case where a large language model (LLM) is used. Such information was suggested by one expert, is used in LLM-related databases (LVEs, ), and is used in model cards (Derczynski et al., 2023). Such information may be determined for individual usage, and may thus be adapted freely.
@@ 4.3 Incident information
General information (3.1-3.4) First, the materialization status is collected, as suggested by two experts and as done in one database (ARVA, ). This information distinguishes a vulnerability found during testing (e.g., red-teaming) from an exploit of a deployed system. We further put forward attempting to find the root cause of the incident and report it. To ensure that an incident is indeed an AI security incident as defined in Definition 2.6, the attacker’s malicious intent should recorded. This was suggested by half of our experts, forms part of several databases (ARVA, ; OECD, ; Walker et al., 2024), and was suggested in prior work (Grosse et al., 2024b). Finally, we recommend recording the attack type. For coding purpose and in this paper, we rely on the categorization by Grosse et al. (2024b). Any other framework, however, may also serve this purpose.
Relevant components (3.5-3.7) An AI security incident is further strongly categorized by the resources used by the attacker to carry out the attack (Apruzzese et al., 2023). We thus recommend identifying and reporting the target asset, e.g. whether the model or the entire system was targeted, pointed out in prior work (Bieringer et al., 2022). In addition, we propose to keep track of hardware in relation to the incident, as suggested by one expert. Finally, access rights to all components of the AI (data, model, queries, etc) should be reported to identify how the attacker possibly entered the system. The latter was suggested by one expert and studied in prior work (Grosse et al., 2024a, 2023).
Defensive measures (3.8 and 3.9) The access patterns also give an understanding how which access, in theory, should not have been possible. Along these lines, there is value in understanding what measures were in place before the incident took place, but did not prevent it. To this end, we advocate reporting existing security measures, as suggested by three experts and the literature in model cards (Hupont et al., 2024; Arnold et al., 2019; Adkins et al., 2022). This is in particular relevant for large language models, where implemented guardrails should be reported (Rebedea et al., 2023). Along the same lines, we suggest also collecting fail-safe measures.
@@ 4.4 Incident implications
To understand the severity of an AI security incident, it is crucial to monitor the damage it causes. Monitoring such effects is part of this brief section with elements 4.1 to 4.4. The information in this section may be difficult to derive depending on the role of the reporter—a possible solution for this problem is to rely on a committee or other help.
As the first element, we propose to monitor whether the incident led to non-compliance with existing regulations, a factor brought up by industry participants in security contexts (Bieringer et al., 2022; Grosse et al., 2023). Secondly, any harm that occurred should be well-documented, as suggested by the CSET AI Harm Taxonomy (Hoffman & Frase, 2023) and applied within the AI Incident Database (McGregor, 2021). While it is recognized that determining (or defining) the extent of harm of an AI incident can be inherently uncertain (Paeth et al., 2024), we advocate that it is still worth attempting to record. A different level of abstraction is to use the CIA definition from Section 2.1 to characterize the impact of the incident.
Finally, and beyond the immediate incident, the goal of our taxonomy is to understand vulnerabilities beyond an individual system. We thus advocate explicitly reporting an assessment of whether the incident could affect other AI systems. This was also suggested by three of our experts and is implemented in the Huntr database already (Huntr, ).
@@ 4.5 Appendix
5 Limitations, implementation, Implications and future research
Having discussed the proposed taxonomy, we now revisit known limitations, how the taxonomy should be implemented to be of most use, and finally implications and open research questions.
@@ 5 Limitations, implementation, Implications and future research
Limitations. As stated in the introduction, this taxonomy is a basis for a discussion within the community, not a final version. As such, it has several limitations. Firstly, it does not cover offensive AI, as we consider this usage orthogonal to the security incidents we aim to capture. Secondly, the system currently does not handle AI systems with multiple AI models or components, or recursive usage of AI. While an extension covering several AI systems is possible by filling questions $2.x$ for all covered AI models, there may be better responses to this shortcoming. In general, the taxonomy will have to be adapted for new use cases and trends in AI. Furthermore, security incidents are very application specific [^10] As an example, consider AlphaFold’s prohibited use policy https://github.com/google-deepmind/alphafold3/blob/main/WEIGHTS_PROHIBITED_USE_POLICY.md, and it may be necessary to adapt the taxonomy for specific applications. Furthermore, our current definition of security also covers incidents with “potential” effects. While this enables to cover red-teaming results or vulnerabilities with unknown statuses of exploitation, in the long run it may be more beneficial to align the definition of security definitions including only incidents with identified harm.
@@
Implementation. Many properties desired by the community, such as ease-of-use, machine readability (Sect. 2.4), and reproducibility (Sect. 4.2) depend on the practical implementation of our taxonomy. For example, the taxonomy is best to use when frameworks for attack and harm categorization are well known to the reporter and downstream audience. We thus suggested that external frameworks in certain corresponding questions (for example, attacks and harm questions), can be utilized (Sect. 4.3). Along these lines, machine readability (Sect. 2.4) depends strongly on whether answers are selected from finite lists (such as multiple-choice or drop-down responses) as opposed to open text fields, or defining the taxonomy keys and values in a more generalizable schema (such as XML, RDFa, or other standards). While we most likely cannot eliminate all open text fields, we strongly suggest minimizing them as much as possible in order to allow ease of analysis. Lastly, it would be good to allow the reproducibility of an incident (Sect. 4.2). However, reproducibility directly contradicts the intellectual property of a model release. In this context, it would be possible to reduce the reporting scheme to contain only information supporting understanding of the incident, and not release model weights, for example. Otherwise, a trusted third party could act as an intermediary.
@@
Implications and future work. We have outlined that the exact reporting taxonomy and implementation strongly influence the overall usefulness and applicability of an AI incident reporting tool. Understanding the requirements of new emerging trends within AI (security), new applications, and, machine readability of AI incident data are open research questions. This extends to the tension between the protection of proprietary AI model information and the learning and generalizability of AI security incidents. The draft we present here is intended as a scaffold to determine, discuss, and apply such changes. Orthogonally, but analogously to other areas of incident reporting, usability studies could help to understand and advance both the usefulness of the collected reports (Marcilly et al., 2019; Mazaheri et al., 2015; Wiele & Rantanen, 2015). Another aspect is to report the identity of the attacker. While similar frameworks such as CVE neither cover the identification or the collection of information on the attacker, in the long run approaches obtaining such information would greatly benefit both research and AI users.
@@ 6 Alternative views
The position and suggested taxonomy in this paper is not the only way to approach AI security incident reporting. While the exact taxonomy, as stated, is open for debate, there are more extreme positions to be taken on this topic.
Position—No reporting is needed. One may argue that due to the absence of many AI security incidents, AI security reporting is not (yet) needed. In contrast, we believe that based on existing reports of AI security incidents (Grosse et al., 2024b; McGregor, 2021; ARVA, ) and the idea of proactive security (Biggio & Roli, 2018), incident collection should rather start too early than too late.
Position—CVE reporting is sufficient. One approach is integrating AI security reports into the existing CVE program. While we have highlighted several reasons against this approach (Sect. 2.5), it could instead be argued that the CVE scheme could be altered to better fit AI security incidents. We have reasoned that CVE is too general and does not consider that data on deployed AI models are often withheld by companies. However, regulatory frameworks might require model providers and deployers to make the specific information related to AI security incidents transparent.
Position—Extending general AI incident reporting. A different approach to the one described above could be using the existing infrastructure (Sect. 2.4) for non-security AI incidents and extending it to security. While we argue that the frameworks are sufficiently different and security incidents merit specific attention, existing infrastructure and AI incident communities could be extended towards security.
Position—A smaller, less bureaucratic approach is needed. There may be justified criticism that the 26 questions (Sect. 4) are too many and a bureaucratic burden for the reporters. While we have tried to argue in favor of reusing existing infrastructure such as model cards (Sect. 4), we would like to emphasize that the here listed elements are not a minimum, but an ideal suggestion and wish list of several parties. We are open to discussing which of the 26 questions could be a minimal set to still fulfill the purpose of different stakeholders.
Position–Taxonomies evolve over time. As reasoned above, security reporting infrastructures have been changing over time to accommodate emerging new methods and tools (Sect. 2.5). It is thus expected that a taxonomy of AI security incidents will evolve in a similar fashion over time, especially our proposed taxonomy. Yet, we argue that this draft is valid as a base of discussion and that eventually, a framework is needed to standardize communication between reporting entities and entities receiving reports as required legally.
@@
The current annual damage of cyber crime, when quantified in money, is 8 trillion USD in 2023 alone. [^11] https://www.statista.com/forecasts/1280009/cost-cybercrime-worldwide Understanding and combating such crime is thus of uttermost importance. While virtually all of this number is currently caused by “classical” security issues (unrelated to AI), the deployment of AI systems across sectors is dramatically increasing and AI is likely to be implicated in far more cybercrime losses; AI security vulnerabilities are rather well established (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). At the same time, little is known about the security incidents that occur in the wild on deployed systems (Apruzzese et al., 2023; Grosse et al., 2024b). In aviation, the rigorous study of incidents has lead to a decrease in casualties over time (Li, 2019). Analogously, knowledge about practical AI security incidents could fuel AI security research leading to research in more practical threat models. We hope with our paper to make a significant step towards a taxonomy for AI security incidents that benefits all stakeholders—the developers and deployers affected, but also research, industry, and governmental entities who seek to analyze and learn from these incidents.
Acknowledgments The authors would like to thank Jochen Stängler (German Federal Office for Information Security, BSI) who has given extensive feedback on both the definitions and the draft of the taxonomy.

### v2

@@ 3 Approach
To derive a taxonomy for reporting and describing AI security incidents, we followed a three-step approach. Firstly, we reached out to twelve industry experts (working in AI red teaming, security industry, AI governance, or market oversight), inquiring about their wish list for a possible reporting scheme. After distilling concepts from these conversations, we matched these concepts across experts, with model card approaches (Section 2.3), and existing databases (Section 2.5). The derived, rudimentary scheme was then refined by iterative discussion rounds among four of the authors. As a consequence, duplicates were removed. For example, some experts raised the question of whether the model was re-used, while others noted the relevance of keeping track of supply-chain-related information regarding the model. Here, the fourth-fold structure of the taxonomy emerged. In this process, we added meta-information from CSET (Hoffmann et al., ) like incident date, annotator, and whether the AI system was applied in a cyber-physical context.
When both, taxonomy draft and definitions, had sufficiently converged, three of the authors used the tentative structure to tag incidents from existing databases. The tagging was overall consistent, except in fields like the classification of harm, where two authors had significantly less background than the third; showing the need for guidelines. Other examples with disagreement related to the generalizeability of incidents, a challenging question for which, to the best of our knowledge, standard assessment procedures are not yet established. Afterwards, the proposal (in the form of this paper) was shared with the initial and additional experts for a broader, final round of feedback. Here, no significant changes to the taxonomy occurred.
@@ 4 AI security incident taxonomy
Our proposed taxonomy consists of five parts, where the last contains only an open text field for further comments and feedback. The parts are, (1) the incident’s meta data, (2) information about the targeted AI system, (3) specifics of the security incident, and lastly (4) implications of the incident. We now revisit the individual questions and references in each part. An example of the final questionnaire is depicted in Table 2-4 in the appendix and Figure 1 gives an abstract overview on our proposed taxonomy.
Figure 1: Abstract template of our proposed taxonomy for reporting and describing AI security incidents.
4.1 Meta data (T1.1-T1.4)
The metadata is a short block containing the incident’s high-level information, and is, under variations, part of many existing databases (Huntr, ; McGregor, 2021; LVEs, ; ARVA, ; OECD, b). This information forms the first section with taxonomy elements T1.1-T1.4.
We opted to assign an incident ID, a name or ID of the incident reporter, their organization, and the alleged date (or interval, or unknown) of the incident. It is important to remark that an AI security incident may be reported anonymously. Although, as stated above, any question of our proposal can be dropped, we want to emphasize that the reporter and their organization may be pseudonymized or kept fully anonymous, depending on the context of usage.
@@ 4.2 Targeted AI system
To understand an AI security incident, it is crucial to describe the AI system that formed part of the incident. Hence, in the second section of our taxonomy, we first collect information about the model, its life cycle stage and use, its origin, and application-specific information. Ideally, this information is directly imported from a model card or similar documentation to reduce required efforts.
@@
Model information (T2.1-T2.3). We first collect basic information about the model. This includes the model type, for example, whether a model is generative or predictive; a specific reply here could be a support vector machine. Such information was suggested by half of our experts and is used in several model card approaches (Mitchell et al., 2019; Arnold et al., 2019; Adkins et al., 2022). We then collect more specific information like the exact version of the model, information also encompassed in many model cards (Mitchell et al., 2019; Adkins et al., 2022) and requested by a fourth of our experts. Lastly, if possible, the model should be specified as detailed as possible (e.g., model weights), to enhance reproducibility, as suggested by another quarter of our experts.
@@
Life-cycle stage and intended use (T2.4-T2.7). In addition to the model within the AI system, understanding the system context is crucial. To this end, we also collect lifecycle information as specified by ISO [^11] ISO/IEC 22989 Information technology — Artificial intelligence — Artificial intelligence concepts and terminology, https://www.iso.org/standard/74296.html, as suggested by two of our experts. As suggested by the CSET AI Harm Framework (Hoffman & Frase, 2023) annotation guide, we also collect the domain where the AI system is deployed (Hoffmann et al., ). Intuitively, a deployment in physical domains has different implications than in the digital space. In addition, we keep track of the intended use of the AI system, loosely following the EU AI Act (Council of European Union, 2024), and as done in many model card approaches (Mitchell et al., 2019; Hupont et al., 2024; Arnold et al., 2019; Adkins et al., 2022). To take into account specific exploits, we also advocate denoting a usage during the incident if different from the intended use.
@@
Model supply chain (T2.8-T2.10). Furthermore, it is crucial to understand possible (external) influences on the AI system stemming from the supply chain, as also suggested by two of our experts and in several AI model card approaches (Mitchell et al., 2019; Gebru et al., 2021; Adkins et al., 2022). Following scientific literature of AI security, we opted to divide the supply chain into three major areas: data (Mir, 2024), model weights (Gu et al., 2019), and code (Grosse et al., 2020; Shumailov et al., 2021). Any of these may allow the import of security vulnerabilities. Especially for data, prior work suggests which information could be collected (Bender & Friedman, 2018).
@@
Application (T2.11). Instances of AI security incidents may need specific details to be reproducible or understandable. As an example, we suggest collecting the system prompt in the case where a large language model (LLM) is used. Such information was suggested by one expert, is collected in LLM-related databases (LVEs, ), and model cards (Derczynski et al., 2023).
@@ 4.3 Incident information
General information (T3.1-T3.4). First, the materialization status is collected, as suggested by two experts and as done in one database (ARVA, ). This information distinguishes a vulnerability found during testing (e.g., red-teaming) from an exploit of a deployed system. We further put forward attempting to find the root cause of the incident and report it. To ensure that an incident is indeed an AI security incident as defined in Definition 2.6, the attacker’s malicious intent should be recorded. This was suggested by half of our experts, forms part of several databases (ARVA, ; OECD, b; Walker et al., 2024), and was suggested in prior work (Grosse et al., 2024b). Finally, we recommend recording the attack type. For our tagging, we relied on the categorization by Grosse et al. (2024b). Any other framework, however, may also serve this purpose.
Relevant components (T3.5-T3.7). An AI security incident is further categorized by the resources used by the attacker to carry out the attack (Apruzzese et al., 2023). We thus recommend identifying and reporting the target asset, e.g. whether the model or the entire system was targeted, pointed out in prior work (Bieringer et al., 2022). In addition, we propose to keep track of used hardware, as suggested by one expert. Finally, access rights to all components of the AI system (data, model, queries, etc) should be reported to identify how the attacker possibly entered the system. The latter was suggested by one expert and studied in prior work (Grosse et al., 2024a, 2023).
Defensive measures (T3.8-T3.9). Access patterns also give an understanding which access may have prevented the attack. Along these lines, there is value in understanding what measures were in place before the incident took place, but did not prevent it. To this end, we advocate reporting existing security measures, as suggested by a quarter of our experts and the literature in model cards (Hupont et al., 2024; Arnold et al., 2019; Adkins et al., 2022). This is in particular relevant for large language models, where implemented guardrails should be reported (Rebedea et al., 2023). Along the same lines, we suggest also collecting fail-safe measures.
@@ 4.4 Incident implications
To understand the severity of an AI security incident, it is crucial to monitor resulting damage. We capture such effects in this brief section with elements T4.1-T4.4. The information in this section may be difficult to derive depending on the role of the reporter—a possible solution for this problem is to rely on a committee or internal risk management teams.
As the first element, we propose to monitor whether the incident led to non-compliance with existing regulations, a factor brought up by industry participants in security contexts (Bieringer et al., 2022; Grosse et al., 2023). Secondly, any harm that occurred should be well-documented, as suggested by the CSET AI Harm Taxonomy (Hoffman & Frase, 2023) applied within the AI Incident Database (McGregor, 2021). While determining (or defining) the extent of harm of an AI incident can be inherently uncertain (Paeth et al., 2024), we advocate that it is still worth attempting to record harms resulting from security incidents.
A different level of abstraction is to use the CIA definition from Section 2.2 to characterize the impact of the incident. Finally, and beyond the immediate incident, the goal of our taxonomy is to understand vulnerabilities beyond an individual system. We thus advocate explicitly reporting an assessment of whether the incident could affect other AI systems. This was also suggested a fourth of our experts and is implemented in the Huntr database already (Huntr, ).
@@ 4.5 Appendix
5 Limitations, implementation, implications and future work
Having discussed our taxonomy, we revisit its limitations or possible extensions. We then discuss desirable features when implementing the taxonomy, for example, machine readability. We conclude with related open questions, for example the tension between intellectual property and the information collected.
@@ 5 Limitations, implementation, implications and future work
Limitations. Our taxonomy is a basis for a discussion within the community. As such, it has several limitations. Firstly, it does not cover offensive AI, in other words using AI to carry out attacks. Such usage is, in our opinion, orthogonal to security incidents. Secondly, the proposal currently does not handle AI agents, multi agent systems, AI systems with multiple AI models or components, or recursive usage of AI. While an extension covering several AI systems is possible by filling questions $2.x$ for all covered AI systems, there may be better solutions to this shortcoming. In general, the taxonomy and underlying definitions will have to be adapted for new use cases and trends in AI. Furthermore, AI and its security incidents are very application specific [^12] As an example, consider AlphaFold’s prohibited use policy https://github.com/google-deepmind/alphafold3/blob/main/WEIGHTS_PROHIBITED_USE_POLICY.md, and it may be necessary to adapt the taxonomy for specific applications. Furthermore, our current definition of security also covers incidents with “potential” effects. While this enables to cover red-teaming results or vulnerabilities with unknown statuses of exploitation, in the long run it may be more beneficial to align the definition of security to include only incidents with identified harm.
@@
Implementation. Many properties desired by the community, such as ease-of-use, machine readability (Section 2.5), and reproducibility (Section 4.2) depend on the practical implementation of our taxonomy. For example, the taxonomy is best to use when frameworks for attack and harm categorization are well known to the reporter and additional stakeholders involved in incident reporting and mitigation. We thus suggested that external and pre-existing frameworks in certain corresponding questions (for example, attacks and harm questions), can be utilized (Section 4.3). Along these lines, machine readability (Section 2.5) depends strongly on whether answers are selected from finite lists (such as multiple-choice or drop-down responses) as opposed to open text fields, or defining the taxonomy keys and values in a more generalizable schema (such as XML, RDFa, or other standards). While we most likely cannot eliminate all open text fields, we strongly suggest minimizing them as much as possible in order to allow ease of analysis. Lastly, it would be good to allow the reproducibility of an incident (Section 4.2). However, reproducibility directly contradicts the intellectual property of a model release. In this context, it would be possible to reduce the reporting scheme to contain only information supporting understanding of the incident, and not release model weights, for example. This would also include enforcing exact versioning to allow referencing a model without sharing it. Otherwise, a trusted third party could act as an intermediary.
@@
Implications and future work. We have outlined that the exact reporting taxonomy and implementation strongly influence the overall usefulness and applicability of an AI incident reporting framework. Understanding the requirements of new emerging trends within AI (security), new applications, and, machine readability of AI incident data are open research questions. This extends to the tension between the protection of proprietary AI model information and the learning and generalizability of AI security incidents. The draft we presented here is intended as a scaffold to determine, discuss, explore and apply these facets.
Orthogonally, but analogous to other areas of incident reporting, usability studies could help to understand and advance both the usefulness of the collected reports (Marcilly et al., 2019; Mazaheri et al., 2015; Wiele & Rantanen, 2015). We have also outlined that in some cases, like supply chain (T2.8-T2.10, Section 4.2), attacks (T3.4, Section 4.3), and harm (T4.2, Section 4.4), a decision has to be taken which frameworks are to be chosen. This is an additional avenue for future work. Another aspect of our taxonomy could be to identify attacker’s identity. While similar frameworks such as CVE neither cover the identification or the collection of information on the attacker, in the long run approaches obtaining such information would greatly benefit both research, AI deployers and stakeholders.
@@ 6 Alternative views
The position and taxonomy proposed in this paper are not the only way to approach AI security incident reporting. While the exact taxonomy, as stated, is open for debate, there are other positions to be taken on this topic.
Position—No reporting is needed. One may argue that due to the absence of many real world AI security incidents, AI security reporting is not (yet) needed. In contrast, we believe that based on initial reports of AI security incidents (Grosse et al., 2024b; McGregor, 2021; ARVA, ) and the idea of proactive security (Biggio & Roli, 2018), appropriate incident collection and analysis is crucial. In other words, incident reporting should start now to enable proactive, as opposed not reactive security treatment.
Position—CVE reporting is sufficient. Another position is to integrate AI security reports into the CVE program(Fazelnia et al., 2024), possibly altering the CVE scheme to better fit AI security incidents. We have highlighted several counterarguments (Section 2.6), including that CVE is too general and does not consider that data and AI models are often withheld by companies. However, regulatory frameworks might require model providers and deployers to make the specific information related to AI security incidents transparent, enabling CVE to be used.
Position—Extending general AI incident reporting. A different approach to the one described above could be relying on existing infrastructure (Section 2.5) for non-security AI incidents and extending it to security. While we argue that the frameworks are sufficiently different and security incidents merit specific attention, existing infrastructure and AI incident communities could be extended to security.
Position—A smaller set of information is needed. There may be criticism that the over 25 questions (Section 4) are too many and a bureaucratic burden for the reporters. While we argue in favor of reusing existing infrastructure such as model cards (Section 4), we would like to emphasize our taxonomy is not a minimal, but an ideal suggestion to likely satisfy the requirements of all affected or involved stakeholder groups. We are open to discussions around and a creation of a minimal subset of our taxonomy.
Position–Taxonomies evolve over time. As reasoned above, security reporting infrastructures have been changing over time to accommodate emerging new methods and tools (Section 2.6). It is thus expected that a taxonomy of AI security incidents will evolve in a similar fashion over time. Yet, we argue that this draft is valid as a base of discussion. Eventually, a maintained framework is needed to standardize communication between reporting entities (e.g., providers of general purpose AI models with systemic risk) and entities receiving reports as required by applicable regulation (e.g., European AI Office).
@@
The current annual damage of cyber crime, when quantified in monetary terms, was eight trillion USD in 2023 alone. [^13] https://www.statista.com/forecasts/1280009/cost-cybercrime-worldwide Understanding and combating such crime is thus of uttermost importance. While the majority of this number is currently caused by “classical” security issues (unrelated to AI), the deployment of AI systems across sectors is dramatically increasing and AI is likely to be implicated in far more cybercrime losses. We hope with our paper to make a significant step towards a taxonomy for AI security incidents that benefits all stakeholders—the developers and deployers affected, but also research, industry, and governmental entities who seek to analyze and learn from these incidents.
Acknowledgments We would like to thank our anonymous experts and their investment in giving initial inputs and feedback on this work. Some of our experts consented to being explicitly named, including In addition, the authors would like to thank Brian Pendleton, whom we thank especially for his feedback regarding CVE/CWE; Hyrum Anderson, Cisco; Leon Derczynski, NVIDIA Corporation, whom we thank for his feedback on our definitions, methodology, and references; and Yuvaraj Govindarajulu, AIShield, Bosch Global Software Technologies. In addition, we would like to thank Sean McGregor, AIID (AI incident database) for his feedback.